OpenStudy (anonymous):

A test car moves at a constant speed around a circular track. If the car is 48.2 m from the track's center and has a centripetal acceleration of 8.05 m/s^2, what is the car's tangential speed?

OpenStudy (anonymous):

How do I go about solving this problem? I know I have to use this equation:\[F _{c} = (m \times v ^{2}) \div r\] Where r equals 48.2 meters and Fc equals 8.05 m/s^2. The answer is 19.7 m/s, but how do I get to that?

OpenStudy (anonymous):

Centripetal accl = V^2 / r v = sqrt ( Centripetal accl*r) Solve and Find the Value of V..